OpenStudy (anonymous):

Express the complex number in trigonometric form. -2 2(cos 90° + i sin 90°) 2(cos 0° + i sin 0°) 2(cos 180° + i sin 180°) 2(cos 270° + i sin 270°)

OpenStudy (badhi):

\(-2 = -2(1) + i(0)\) \(-2 = -2(\cos\theta)+i(\sin\theta)\) \[\cos \theta = 1 , \sin \theta = 0\] Find the relavant angle - \(\theta\)

OpenStudy (anonymous):

I don't know how to ....

OpenStudy (anonymous):